OpenStudy (anonymous):

The dimensions of a rectangular prism are the same as a Pythagorean triangle with legs of 8 and 15. The volume of the prism is just give formula

Directrix (directrix):

Have you found the third side of the right triangle yet? I'm going to upload a list of Pythagorean triples, not that the attachment is needed.

Directrix (directrix):

Some people memorize a few of the Pythagorean triples to use in working problems. They are based on the Pythagorean theorem so if you know the P.T., you can still find the third length of this prism.

OpenStudy (anonymous):

ok il figure it out thank you

Directrix (directrix):

8^2 + 15^2 = h^2 17 = h See attachment.

Directrix (directrix):

Questions?

Directrix (directrix):

A "Pythagorean Triple" is a set of positive integers, a, b and c that fits the rule: a^2 + b^2 = c^2 And when you make a triangle with sides a, b and c it will be a right angled triangle (see Pythagoras' Theorem for more details).

OpenStudy (anonymous):

8 * 15 * 17 = 2040 units³ thanks